OpenStudy (anonymous):

Boris is kayaking in a river with a 6 mph current. Suppose that he can kayak 4 miles upstream in the same amount of time as it takes him to kayak 9 miles downstream. Find the speed (mph) of Boris’s kayak in still water.

OpenStudy (anonymous):

Just cant set it up, i got like a 4(x-6)=1 and a 9(x+6) = 1?

OpenStudy (anonymous):

v t = S Up stream: ( v - 6 ) t = 4 miles Down stream : ( v + 6) t = 9 miles